# Diverging entanglement of critical magnons in easy-axis antiferromagnets

### v1

@@ I Introduction
In this work, we theoretically study the physics of magnons in general antiferromagnetic materials in the presence of easy-axis anisotropy and an external magnetic field, which are oriented perpendicular to each other [Fig. 1(b)]. The interplay between the two-mode and single-mode squeezing effects due to the antiferromagnetic interaction and the anisotropy, respectively, drives the magnetic phase transition among the spin orderings, controlled by an external magnetic field. Notably, multiple magnetic phase transitions emerge in the ferrimagnets. We find that the magnon-magnon entanglement diverges at all critical points, demonstrating that the critical antiferromagnetic magnon entanglement induces the magnetic instability. Having established the theory of magnons in anisotropic antiferromagnets, we consider a cavity magnonics setup where the antiferromagnetic magnons are coupled to quantized cavity photons. We show that the spontaneous magnetization of the ferromagnets accompanies the cavity photon condensation, thereby translating the magnetic phase transitions due to the critical magnon entanglement to the superradiant phase transition [30, 31, 26]. Moreover, both the quadrature squeezing and the number of photons diverge at the critical point. Measurement of cavity photons therefore provides information about magnonic entanglement and magnetic criticality in these systems.
@@ III.1 Mean-field solutions
In the latter solution, $\theta_{\sigma}$ takes two distinct values, with $\theta_{A}$ for one sublattice and $\theta_{B}$ for the other. The former (latter) solution is stable if the right-hand side of the above equation is larger (lesser) than one. Therefore, there is a critical field strength $B_{0}^{c}\equiv(2JSn_{1}+2KS)/\gamma$ that separates two distinct phases. Note that the spin Hamiltonian possesses the mirror symmetry along the $x$ axis, $\mathcal{M}_{y}$, whose symmetry operation is given by
@@ III.2 Critical entanglement of magnons
where $\omega^{\sigma}_{0}=-JS_{\bar{\sigma}}n_{1}\cos(\theta_{A}-\theta_{B})+K_{\sigma}S_{\sigma}(2-3\sin^{2}\theta_{\sigma})+\gamma B_{0}\sin\theta_{\sigma}$, $\omega^{\sigma}_{s}=(K_{\sigma}S_{\sigma}/2)\sin^{2}\theta_{\sigma}$, and $h^{\pm}_{\mathbf{k}}=(J\sqrt{S_{A}S_{B}}n_{1}\gamma_{\mathbf{k}}/2)(1\pm\cos(\theta_{A}-\theta_{B}))$. Here, $\theta_{\sigma}$ represents the mean-field solution and is a function of system parameters, including the magnetic field strength. We note that there are several types of squeezing terms. First, there are squeezing processes among magnons within a single sublattice, whose magnitude is given by $\omega_{s}^{\sigma}$. They are single-mode and two-mode squeezing for the Kittel (${\bf k}=0$) mode and finite momentum modes, respectively. In addition, there are two-mode squeezing terms between different sublattice magnons proportional to $h^{-}_{\bf k}$. In addition, there is a beam-splitter-type magnon-hopping term between the sublattice magnons with a hopping strength $h^{+}_{\bf k}$. These single-mode and two-mode squeezing terms, together with the beam-splitter type hopping interaction, offer an important toolbox to engineer magnon states and magnon-magnon coupling in magnetic materials [8, 9, 10]. The role of the single-mode squeezing term in modifying the spin angular momentum of the magnon eigenmode [37, 22, 26] and the criticality of the ferromagnet [26] has been investigated. As we demonstrate below, the interplay between various inter- and intra-sublattice squeezing processes drives the phase transition among the spin orderings investigated previously, through the critical entanglement of magnons.
@@
Let us now investigate the correlation between the different sublattice magnons by calculating the entanglement. The various squeezing terms and hopping terms in the magnon Hamiltonian [Eq. (11)] contribute to the entanglement, unlike the case where the two-mode squeezing parameter entirely represented the entanglement between sublattices [23, 35]. To measure the entanglement, we calculate and plot the von Neumann entropy $E_{ab}$ between the uniform magnon modes $a=a_{\mathbf{k}=0}$ and $b=b_{\mathbf{k}=0}$ of different sublattices in Fig. 3 using the covariance matrix [39]. The entanglement is nonzero in the absence of the magnetic field due to two-mode squeezing from antiferromagnetic exchange interaction. As the magnetic field increases, the entanglement diverges at all critical points, indicating the divergence of the correlation, similar to the divergence of single-mode squeezing in the easy-axis ferromagnet [26]. Note that diverging entanglement appears finite in the figure due to numerical precision. It is interesting to observe that the two-mode squeezing terms proportional to $h^{-}_{\bf k}$ become zero at the critical point; therefore, the divergence of the entanglement is due to the interplay between the single-mode squeezing and the number-conserving terms. Our analysis shows that the anisotropy and the perpendicular magnetic field can induce arbitrarily large entanglement entropy of magnons, facilitating the engineering of magnon-magnon entanglement [25, 40, 41].
@@
We also plot the number of magnons $\langle a^{\dagger}a\rangle$ and $\langle b^{\dagger}b\rangle$ in the ground state as a function of the magnetic field strength in Fig. 3. These values are nonzero across all ranges due to the squeezing. Unlike the entropy, the magnon numbers are distinct for each component, with the difference proportional to the asymmetry in anisotropy strength and spin amplitude. The magnon numbers also diverge at all critical points as the infinite bipartite entropy between two bosonic modes requires an infinite fluctuation in the number of excitations. We also find an interesting non-monotonic behavior of the magnon number; namely, the magnon number is strongly suppressed before this divergence at the phase transition (see arrows in Fig. 3(c) and (e)). At this point, the coefficients of the squeezing terms $\omega^{\sigma}_{s}$ and $h^{-}_{\bf k}$ become nearly identical. As the magnetic field increases, both angles $\theta_{\sigma}$ approach $\pi/2$, leading to a situation where $\omega^{\sigma}_{s}$ and $h^{-}_{\bf k}$ become vanishingly small, but remain finite unless $J\gg K_{\sigma}$. As a result, the coefficients of all squeezing terms have nearly identical small amplitudes, leading to the highly suppressed magnon number in the ground state. Despite this suppression, the ground state never becomes an exact vacuum state under any parameter conditions. Both the magnon number suppression and divergence are experimentally observable consequences of the competing magnon squeezing effects.
@@ IV Cavity magnonics
We consider an antiferromagnet or a ferrimagnet, described by Eq. (1), placed within a cavity where the cavity magnetic field is oriented along the $z$ direction. We focus solely on the uniform magnon mode that couples to the uniform cavity field under the dipole approximation. Additionally, for simplicity, we assume symmetric bare spin-cavity coupling strengths for both sublattices. The Hamiltonian for the combined system of the cavity and the magnet is given by
$H^{\prime}_{0}=H_{0}+\omega_{c}c^{\dagger}c-g(c+c^{\dagger})\sum_{i,\sigma}\sqrt{\frac{2}{S_{\sigma}N}}S^{z}_{\mathbf{r}_{i}},$ (12)
@@
$\mathcal{E}_{\rm cm}=$ $\mathcal{E}_{\rm m}-2gc_{0}\sum_{\sigma}\sqrt{\frac{2S_{\sigma}}{N}}\cos\theta_{\sigma}+\omega_{c}c^{2}_{0},$ (13)
@@
$H^{\text{cm}}=$ $\frac{\omega^{A}_{0}}{2}a^{\dagger}a+\frac{\omega^{B}_{0}}{2}b^{\dagger}b+\frac{\omega_{c}}{2}\tilde{c}^{\dagger}\tilde{c}+\omega^{A}_{s}a^{2}$ (14)
@@
where $\tilde{c}=c-c_{0}$ is the effective cavity photon operator with the photon condensation
@@
$a=a_{\mathbf{k}=0}$ and $b=b_{\mathbf{k}=0}$ are magnon operators of uniform modes, $h^{\pm}=h^{\pm}_{\mathbf{k}=0}$, and $\tilde{g}_{\sigma}=g\sin\theta_{\sigma}$. We first note that photon condensation occurs with non-zero order parameter $c_{0}$ only when the net angular momentum exists (green lines in Fig. 2), indicating the emergence of the superradiant phase. We solve the cavity magnonics Hamiltonian given in Eq. (14), which leads to the hybridization of cavity photons and magnons into a magnon-polariton. We plot the frequency dispersion of the magnon-polariton for both antiferromagnetic and ferrimagnetic regimes in Fig. 4(a) and (b). The avoided crossings in the dispersion reveal the magnon-photon interaction despite their small size in the plot. Notably, the low-lying magnon-polariton softens and becomes critical at the superradiant phase transition point with a power-law, $\omega\propto|B_{0}-B^{c}_{0}|^{1/2}$. We also investigate the quantum fluctuation of the cavity field by calculating the variance of quadrature variables defined as follows,
@@
which can be measured using homodyne detection inside the cavity. We plot the variance of quadratures $\langle(\Delta x_{c})^{2}\rangle$ and $\langle(\Delta p_{c})^{2}\rangle$, along with the number of cavity photons $\langle c^{\dagger}c\rangle$ in Figs. 4(c) and (d). Both the variance of the quadrature of $x_{c}$ and the cavity photon number diverge at all critical points, an important characteristic of a superradiant phase transition. Furthermore, the critical field strengths of all critical points are influenced by the magnon-photon coupling. As illustrated in Fig. 4(e) and (f), the field strength at which the photon number diverges shifts upwards as the coupling strength increases.

### v2

@@ I Introduction
In this work, we theoretically study the physics of magnons in general antiferromagnetic materials in the presence of easy-axis anisotropy and an external magnetic field, which are oriented perpendicular to each other [Fig. 1(b)]. The interplay between the two-mode and single-mode squeezing effects due to the antiferromagnetic interaction and the anisotropy, respectively, drives the magnetic phase transition among the spin orderings, controlled by an external magnetic field. Notably, multiple magnetic phase transitions emerge in the ferrimagnets. We find that the magnon-magnon entanglement diverges at all critical points, demonstrating that the magnetic instability induces the critical antiferromagnetic magnon entanglement. Having established the theory of magnons in anisotropic antiferromagnets, we consider a cavity magnonics setup where the antiferromagnetic magnons are coupled to quantized cavity photons. We show that the spontaneous magnetization of the ferromagnets accompanies the cavity photon condensation, thereby translating the magnetic phase transitions due to the critical magnon entanglement to the superradiant phase transition [30, 31, 26]. Moreover, both the quadrature squeezing and the number of photons diverge at the critical point. Measurement of cavity photons therefore provides information about magnonic entanglement and magnetic criticality in these systems.
@@ III.1 Mean-field solutions
In the latter solution, $\theta_{\sigma}$ takes two distinct values, with $\theta_{A}$ for one sublattice and $\theta_{B}$ for the other. The former (latter) solution is stable if the right-hand side of the above equation is larger (lesser) than one. Therefore, there is a critical field strength $B_{0}^{c}\equiv(2JSn_{1}+2KS)/\gamma$ that separates two distinct phases. Note that the spin Hamiltonian possesses the mirror symmetry along the $y$ axis, $\mathcal{M}_{y}$, whose symmetry operation is given by
@@ III.2 Critical entanglement of magnons
where $\omega^{\sigma}_{0}=-JS_{\bar{\sigma}}n_{1}\cos(\theta_{A}-\theta_{B})+K_{\sigma}S_{\sigma}(2-3\sin^{2}\theta_{\sigma})+\gamma B_{0}\sin\theta_{\sigma}$, $\omega^{\sigma}_{s}=(K_{\sigma}S_{\sigma}/2)\sin^{2}\theta_{\sigma}$, and $h^{\pm}_{\mathbf{k}}=(J\sqrt{S_{A}S_{B}}n_{1}\gamma_{\mathbf{k}}/2)(1\pm\cos(\theta_{A}-\theta_{B}))$. Here, $\theta_{\sigma}$ represents the mean-field solution and is a function of system parameters, including the magnetic field strength. We neglected the higher-order magnon terms as these terms are typically suppressed by the size of the magnet; see Appendix A. We note that there are several types of squeezing terms. First, there are squeezing processes among magnons within a single sublattice, whose magnitude is given by $\omega_{s}^{\sigma}$. They are single-mode and two-mode squeezing for the Kittel (${\bf k}=0$) mode and finite momentum modes, respectively. In addition, there are two-mode squeezing terms between different sublattice magnons proportional to $h^{-}_{\bf k}$. In addition, there is a beam-splitter-type magnon-hopping term between the sublattice magnons with a hopping strength $h^{+}_{\bf k}$. These single-mode and two-mode squeezing terms, together with the beam-splitter type hopping interaction, offer an important toolbox to engineer magnon states and magnon-magnon coupling in magnetic materials [8, 9, 10]. The coefficients of these squeezing terms as a function of the magnetic field are shown in Fig. 4, which show the relative significance of each term. The role of the single-mode squeezing term in modifying the spin angular momentum of the magnon eigenmode [37, 22, 26] and the criticality of the ferromagnet [26] has been investigated. As we demonstrate below, the interplay between various inter- and intra-sublattice squeezing processes drives the phase transition among the spin orderings investigated previously, through the critical entanglement of magnons.
@@
Let us now investigate the correlation between the different sublattice magnons by calculating the entanglement. The various squeezing terms and hopping terms in the magnon Hamiltonian [Eq. (11)] contribute to the entanglement, unlike the case where the two-mode squeezing parameter entirely represented the entanglement between sublattices [23, 35]. Here, we focus on the entanglement property of the uniform modes, which becomes critical, and we present the analysis of non-uniform modes with finite momentum in Appendix B. To measure the entanglement, we calculate and plot the von Neumann entropy $E_{ab}$ between the uniform magnon modes $a=a_{\mathbf{k}=0}$ and $b=b_{\mathbf{k}=0}$ of different sublattices in Fig. 3 using the covariance matrix [39]. The entanglement is nonzero in the absence of the magnetic field due to two-mode squeezing from antiferromagnetic exchange interaction. As the magnetic field increases, the entanglement diverges at all critical points, indicating the divergence of the correlation, similar to the divergence of single-mode squeezing in the easy-axis ferromagnet [26]. Note that diverging entanglement appears finite in the figure due to numerical precision. It is interesting to observe that the two-mode squeezing terms proportional to $h^{-}_{\bf k}$ become zero at the critical point; therefore, the divergence of the entanglement is due to the interplay between the single-mode squeezing and the number-conserving terms. Our analysis shows that the anisotropy and the perpendicular magnetic field can induce arbitrarily large entanglement entropy of magnons, facilitating the engineering of magnon-magnon entanglement [25, 40, 41].
@@
We also plot the number of magnons $\langle a^{\dagger}a\rangle$ and $\langle b^{\dagger}b\rangle$ in the ground state as a function of the magnetic field strength in Fig. 3. These values are nonzero across all ranges due to the squeezing. Unlike the entropy, the magnon numbers are distinct for each component, with the difference proportional to the asymmetry in anisotropy strength and spin amplitude. The magnon numbers also diverge at all critical points as the infinite bipartite entropy between two bosonic modes requires an infinite fluctuation in the number of excitations. We also find an interesting non-monotonic behavior of the magnon number; namely, the magnon number is strongly suppressed before this divergence at the phase transition (see arrows in Fig. 3(c) and (e)). At this point, the coefficients of the squeezing terms $\omega^{\sigma}_{s}$ and $h^{-}_{\bf k}$ become nearly identical. As the magnetic field increases, both angles $\theta_{\sigma}$ approach $\pi/2$, leading to a situation where $\omega^{\sigma}_{s}$ and $h^{-}_{\bf k}$ become vanishingly small, but remain finite if $J\gg K_{\sigma}$. As a result, the coefficients of all squeezing terms have nearly identical small amplitudes, leading to the highly suppressed magnon number in the ground state. Despite this suppression, the ground state never becomes an exact vacuum state under any parameter conditions. Both the magnon number suppression and divergence are experimentally observable consequences of the competing magnon squeezing effects.
@@ IV Cavity magnonics
We consider an antiferromagnet or a ferrimagnet, described by Eq. (1), placed within a cavity where the cavity magnetic field is oriented along the $z$ direction. As we shall see, this choice of the cavity field direction ensures that the magnetic instabilities induce the critical phenomena of the cavity such as the photon condensation and diverging photon fluctuation. For other choices of the cavity field directions, we present their analysis in Appendix C. We focus solely on the uniform magnon mode, which couples to the uniform cavity field, assuming a small magnet size relative to the wavelength of the cavity field to ensure the validity of the dipole approximation. Additionally, for simplicity, we assume symmetric bare spin-cavity coupling strengths for both sublattices. The Hamiltonian for the combined system of the cavity and the magnet is given by
$H^{\prime}_{0}=H_{0}+\omega_{c}c^{\dagger}c-g(c+c^{\dagger})\sum_{i,\sigma}\sqrt{\frac{2}{S_{\sigma}N}}S^{z}_{\sigma,\mathbf{r}_{i}},$ (12)
@@
$\mathcal{E}_{\rm cm}=$ $\mathcal{E}_{\rm m}-2gc_{0}\sum_{\sigma}\sqrt{\frac{2S_{\sigma}}{N}}\cos\theta_{\sigma}+\frac{\omega_{c}}{N}c^{2}_{0},$ (13)
@@
Figure 4: Coefficients of the magnon Hamiltonian [Eq. (11)] as a function of the magnetic field strength $B_{0}$ in (a) the antiferromagnetic and (b) the ferrimagnetic regimes. Here, $h^{\pm}=h^{\pm}_{\mathbf{k}=0}$. Effective coupling strengths $\tilde{g}_{A,B}$ of sublattices $A$ and $B$ as a function of the magnetic field strength $B_{0}$ in (c) the antiferromagnetic and (d) the ferrimagnetic regimes.
@@
$H^{\text{cm}}=$ $\frac{\omega^{A}_{m}}{2}a^{\dagger}a+\frac{\omega^{B}_{m}}{2}b^{\dagger}b+\frac{\omega_{c}}{2}\tilde{c}^{\dagger}\tilde{c}+\omega^{A}_{s}a^{2}$ (14)
@@
where
$\omega^{\sigma}_{m}=\omega^{\sigma}_{0}+\frac{4g^{2}}{\omega_{c}\sqrt{S_{\sigma}}}\cos\theta_{\sigma}\sum_{\sigma^{\prime}}\sqrt{S_{\sigma^{\prime}}}\cos\theta_{\sigma^{\prime}}$ (15)
is the renormalized magnon frequency, $\tilde{c}=c-c_{0}$ is the effective cavity photon operator with the photon condensation
@@
$a=a_{\mathbf{k}=0}$ and $b=b_{\mathbf{k}=0}$ are magnon operators of uniform modes, $h^{\pm}=h^{\pm}_{\mathbf{k}=0}$, and $\tilde{g}_{\sigma}=g\sin\theta_{\sigma}$. The effective magnon-photon coupling strength $\tilde{g}_{\sigma}$ depends on the magnetic field, reaching its maximum amplitude in the normal phase where the spins are aligned either parallel or antiparallel to each other (Fig. 4). We also note that photon condensation occurs with non-zero order parameter $c_{0}$ only when the net angular momentum exists (green lines in Fig. 2), indicating the emergence of the superradiant phase. We diagonalize the cavity magnonics Hamiltonian given in Eq. (14), which leads to the hybridization of cavity photons and magnons into a magnon-polariton. We plot the frequency dispersion of the magnon-polariton for both antiferromagnetic and ferrimagnetic regimes in Fig. 5(a) and (b). The avoided crossings in the dispersion reveal the magnon-photon interaction despite their small size in the plot. Notably, the low-lying magnon-polariton softens and becomes critical at the superradiant phase transition point with a power-law with an exponent $1/2$, $\omega\propto|B_{0}-B^{c}_{0}|^{1/2}$. We also investigate the quantum fluctuation of the cavity field by calculating the variance of quadrature variables defined as follows,
@@
which can be observed using cavity photon measurements. We plot the variance of quadratures $\langle(\Delta x_{c})^{2}\rangle$ and $\langle(\Delta p_{c})^{2}\rangle$, along with the number of cavity photons $\langle c^{\dagger}c\rangle$ in Figs. 5(c) and (d). Both $\langle(\Delta x_{c})^{2}\rangle$ and $\langle c^{\dagger}c\rangle$ diverge at all critical points with a power-law with an exponent $-1/2$, an important characteristic of a superradiant phase transition [48, 49]. We note that the variance of $p_{c}$, while being below the minimum uncertainty value of $1/2$ due to squeezing, remains finite. Consequently, the product of variances, $\langle(\Delta x_{c})^{2}\rangle\langle(\Delta p_{c})^{2}\rangle$, which quantifies the uncertainty of the cavity field, diverges at the critical point with the same exponent. Furthermore, the critical field strengths of all critical points are influenced by the magnon-photon coupling. As illustrated in Figs. 5(e) and (f), the field strength at which the photon number diverges shifts upwards as the coupling strength increases.
@@ Acknowledgements.
Appendix A Higher-Order Magnon Terms
In the previous sections, we considered only the quadratic terms in the magnon Hamiltonian, neglecting higher-order terms under the assumption that they are effectively suppressed by the size of the magnet. However, near the critical point, where the magnon number increases significantly, it is important to examine the role of these higher-order terms, particularly in the diverging entanglement. To investigate this, we extend the Holstein-Primakoff transformation to include the next-leading-order terms, as described below.
$S^{+}_{\sigma,\mathbf{r}}$ $\simeq\sqrt{2S_{\sigma}}m_{\sigma,\mathbf{r}}-\frac{1}{2\sqrt{2S_{\sigma}}}m^{\dagger}_{\sigma,\mathbf{r}}m^{2}_{\sigma,\mathbf{r}},$ (18)
$S^{-}_{\sigma,\mathbf{r}}$ $\simeq\sqrt{2S_{\sigma}}m^{\dagger}_{\sigma,\mathbf{r}}-\frac{1}{2\sqrt{2S_{\sigma}}}m^{\dagger 2}_{\sigma,\mathbf{r}}m_{\sigma,\mathbf{r}},$
$S^{z}_{\sigma,\mathbf{r}}$ $\simeq S_{\sigma}-m^{\dagger}_{\sigma,\mathbf{r}}m_{\sigma,\mathbf{r}},$
where $\sigma=A,B$ denotes the sublattice, $m_{A,\mathbf{r}}=a_{\bf r}$ and $m_{B,\mathbf{r}}=b_{\bf r}$ represent the magnon annihilation operators.
Figure 6: (a) The quadrature of the magnon along the $p$ direction and (b) the magnon-magnon entanglement between sublattices $A$ and $B$ are plotted along the magnetic field strength $B_{0}$ near the critical point. The insets denote the strength of the higher-order contribution of the magnon.
Using the extended Holstein-Primakoff transformation outlined above, we expanded the spin Hamiltonian [Eq. (1)], which includes the exchange interaction, easy-axis anisotropy, and Zeeman interaction. This expansion introduces third-order terms involving three magnon operators as the next-leading-order contributions as follows,
$H^{(3)}_{\rm h}=H^{(3)}_{J}+H^{(3)}_{K}+H^{(3)}_{B},$ (19)
where
$H^{(3)}_{J}$ $=i\frac{Jn_{1}}{\sqrt{N}}\text{cssc}_{\theta}\Big{(}\sqrt{\frac{S_{A}}{2}}ab^{\dagger}b-\sqrt{\frac{S_{B}}{2}}a^{\dagger}ab$ (20)
[…]
$H^{(3)}_{K}=\frac{i}{\sqrt{N}}\sum_{\sigma}$ $K_{\sigma}\sin\theta_{\sigma}\cos\theta_{\sigma}\sqrt{\frac{S_{\sigma}}{2}}\Big{(}m_{\sigma}m^{\dagger}_{\sigma}m_{\sigma}$ (21)
[…]
$\text{cssc}_{\theta}=\cos\theta_{A}\sin\theta_{B}-\sin\theta\cos\theta$, $\sigma=A,B$ denotes the sublattice, $m_{A}=a$ and $m_{B}=b$ represent the annihilation operators of the uniform magnon mode. For simplicity, we focus on the uniform mode although the higher-order terms generally include non-uniform modes with finite momentum. In the normal phase, these third-order terms vanish due to the mirror symmetry so we additionally calculate the fourth-order terms which are next-leading-order contributions in the normal phase as follows,
$H^{(4)}_{\rm h}=H^{(4)}_{J}+H^{(4)}_{K},$ (23)
where
$H^{(4)}_{J}$ $=\frac{Jn_{1}}{8N}\sqrt{\frac{S_{A}}{S_{B}}}(c^{-}_{\theta}a-c^{+}_{\theta}a^{\dagger})b^{\dagger}b^{2}$ (24)
$+\frac{Jn_{1}}{8N}\sqrt{\frac{S_{B}}{S_{A}}}a^{\dagger}a^{2}(c^{-}_{\theta}b-c^{+}_{\theta}b^{\dagger})$
[…]
$H^{(4)}_{K}=\frac{1}{8N}\sum_{\sigma}$ $K_{\sigma}\sin^{2}\theta_{\sigma}(m_{\sigma}-m^{\dagger}_{\sigma})$ (25)
[…]
and $c^{\pm}_{\theta}=\cos(\theta_{A}-\theta_{B})\pm 1$.
We investigate the contribution of higher-order terms in $H^{(3)}_{\rm h}+H^{(4)}_{\rm h}$ to the critical behavior by numerically computing the quadrature variance along the $p$ direction and the magnon-magnon entanglement near the critical point using the QuTiP package [50]. In the antiferromagnetic regime, we treat the factor $1/N$ as a control parameter that modulates the strength of the quartic term. For the magnon at sublattice $A$, we define the quadrature variable $p=-i(a-a^{\dagger})/\sqrt{2}$ and calculate its variance $\langle(\Delta p)^{2}\rangle=\langle(p-\langle p\rangle)^{2}\rangle$. The results are shown in Fig. 6. We observe that the divergence of both quantities is mitigated upon the introduction of the higher-order terms, which instead leads to an enhancement of these quantities, with a sharp peak at the critical point. Thus, we conclude that the higher-order terms regularize the diverging critical behavior; however, further investigation, including the consideration of finite momentum, is required for a more precise understanding.
Figure 7: (a) The frequency of the magnon $\omega_{l}$ corresponding to the lower dispersion branch and (b) the magnon-magnon entanglement between sublattices $A$ and $B$ are plotted for the antiferromagnetic regime along the magnetic field strength $B_{0}$. Here, $k_{y}$ and $k_{z}$ are fixed to zero, reducing the analysis to variations along $k_{x}$.
Appendix B Magnon Modes with Finite Momentum
In Sec. III, we discussed the magnon-magnon entanglement between sublattices, focusing exclusively on the uniform magnon modes at the $\Gamma$ point. This focus is justified by the fact that these uniform modes drive the magnetic phase transition by closing their frequency gap at the critical magnetic field condition. However, it is natural to ask about the entanglement properties of other non-uniform modes with finite momentum near the critical point.
To explore this, we calculate the magnon frequency and magnon-magnon entanglement for finite momentum $\mathbf{k}$ in Eq. (11). For simplicity, we consider the antiferromagnetic regime and restrict the finite momentum to the $x$ direction. The results, shown in Fig. 7, reveal that for non-uniform modes, the frequency gap is reduced but does not close at the critical point. Similarly, the entanglement is enhanced near the critical point but does not diverge to infinity. Despite this, the sharp reduction in the frequency gap and the pronounced enhancement of entanglement make the phase transition observable even in the non-uniform modes.
Figure 8: Quadratures of the photon canonical operators and photon number in the antiferromagnetic regime for cavity field orientations along (a) the $x$ direction and (b) the $y$ direction.
Appendix C Cavity Fields in Alternative Directions
In Sec. IV, we assumed the cavity field was aligned along the $z$ axis, parallel to the easy-axis anisotropy. However, in general, the cavity field can be oriented in arbitrary directions. Here, we discuss cases where the cavity field is aligned along the $x$ axis or the $y$ axis. For both orientations, the cavity-spin Hamiltonian takes the following form,
$H^{\prime}_{0}=H_{0}+\omega_{c}c^{\dagger}c-g(c+c^{\dagger})\sum_{i,\sigma}\sqrt{\frac{2}{S_{\sigma}N}}S^{x\>\text{or}\>y}_{\sigma,\mathbf{r}_{i}},$ (26)
where $H_{0}$ is the spin Hamiltonian defined in Eq. (1), and other parameters are consistent with the previous sections.
Regardless of the cavity field direction, one can generally expect instability in the Hamiltonian due to the spin-cavity coupling. In particular, when the cavity field is aligned along the $x$ axis, additional instability may arise, potentially causing the spins to rotate toward the $x$ axis if the coupling is sufficiently strong. However, we do not consider these additional instabilities here. Instead, we focus on verifying whether the magnetic instability of the spin Hamiltonian, as studied in previous sections, propagates to the criticality of the cavity photon, leading to photon condensation and diverging cavity fluctuation.
By performing a mean-field analysis, one can derive the cavity magnonics Hamiltonian based on the classical spin configuration and photon condensation. The photon condensation is given by,
$c_{0}=0\>\text{or}\>\frac{g}{\omega_{c}}\sum_{\sigma}\sqrt{2NS_{\sigma}}\sin\theta_{\sigma},$ (27)
depending on the cavity field orientation. Notably, photon condensation does not occur in the $x$-axis case. The total cavity magnonics Hamiltonian is expressed as
$H^{\rm cm}_{x\>\text{or}\>y}=H^{\rm m}_{\mathbf{k}=0}+\omega_{c}\tilde{c}^{\dagger}\tilde{c}+H^{\prime}_{\rm int},$ (28)
where the cavity-spin interaction term is given by,
$H^{\prime}_{\rm int}$ $=-g(\tilde{c}+\tilde{c}^{\dagger})\sum_{\sigma}(m_{\sigma}+m^{\dagger}_{\sigma}),$ (29)
or
$H^{\prime}_{\rm int}=$ $ig(\tilde{c}+\tilde{c}^{\dagger})\sum_{\sigma}\cos\theta_{\sigma}(m_{\sigma}-m^{\dagger}_{\sigma})$ (30)
$+2gc_{0}\sum_{\sigma}\sqrt{\frac{2}{S_{\sigma}N}}\sin\theta_{\sigma}m^{\dagger}_{\sigma}m_{\sigma},$
depending on the cavity field orientation. Here, $\sigma=A,B$ denotes the sublattice, and $m_{A}=a$ and $m_{B}=b$ represent the magnon annihilation operators. The magnon-photon coupling strength remains invariant under all parameters of the spin Hamiltonian in the case of the $x$ axis. For the $y$-axis case, the magnon-photon coupling is absent in the normal phase, although photon condensation causes a shift in the magnon frequency.
For simplicity, we consider the antiferromagnetic regime. We diagonalize the cavity magnonics Hamiltonian and present the photon number and the quadratures of the photon canonical operators in Fig. 8. Unlike the case where the cavity field is aligned along the $z$ axis, these quantities do not diverge at the critical point where the magnon-polariton frequency vanishes. Therefore, we conclude that aligning the cavity field along the $z$ axis is essential for observing the magnetic phase transition through cavity photon measurements.